OpenStudy (anonymous):

-2a squared if a = -5/2

OpenStudy (anonymous):

(-2)(-5/2)^2 =-2*25/4 =-25/2

OpenStudy (anonymous):

what if the options r these 25 -25

OpenStudy (anonymous):

It's 25. (-2*-5/2)^2 =(-5)^2 =25

OpenStudy (anonymous):

-121/2 and 12 1/2

OpenStudy (anonymous):

what should b done?

OpenStudy (anonymous):

u want me to add both f them?